If, in the reaction below, 83.7 grams of Fe producers 150.0 grams of Fe2S3 what is the % yield?
1
Expert's answer
2020-05-12T11:35:50-0400

83.7g________________________x

2Fe____=>_____Fe2S3

2*55.85 g/mol_________(2*55.85+3*32.06)g/mol = 207.88 g/mol


In case of 100% yield reaction produces:

x = m (Fe2S3) = 83.7 * 207.88 / (2*55.85) = 155.77 g, so the yield is:


w = (150/155.77) * 100% = 96.3%